OpenStudy (anonymous):

Describe the pattern 2,6,12,20,30 Describe the pattern 2,6,12,20,30 @Mathematics

OpenStudy (asnaseer):

work out the difference between each term - that should give you a clue...

OpenStudy (anonymous):

I understand that , ive been having some trouble with sequences in class.

OpenStudy (asnaseer):

so what do you get when you work out the difference between each term in this case?

OpenStudy (asnaseer):

where are you getting stuck?

OpenStudy (anonymous):

determining the pattern, the and finding the rule for the nth term

OpenStudy (asnaseer):

you need to break this up into steps. 1. first write down the difference between each pair of terms on a piece of paper. 2. then look at this new sequence and see if you can spot a pattern. 3. lastly try and use this to work out the equation for the nth term. so, why don't you start with steps 1 and 2 - what do you get?

OpenStudy (asnaseer):

have you learnt about arithmetic progressions/sequences?

OpenStudy (anonymous):

the pattern looks like 4,6,8, and 10

OpenStudy (asnaseer):

ok - so now you know the difference between each term is an increasing sequence of even numbers

OpenStudy (asnaseer):

so, we can write the sequence like this: 2, 2+4, 2+4+6, 2+4+6+8, 2+4+6+8+10 does that help?

OpenStudy (anonymous):

yeah so

OpenStudy (asnaseer):

each term is the sum of an arithmetic progression. do you know the formulae for the sum of an arithmetic progression?

OpenStudy (anonymous):

is it an=a1+ (n-1)d ?

OpenStudy (asnaseer):

yes, so the 1st term is just the sum up to n=1 the 2nd term is the sum up to n=2 etc do you think you can work this out now?

OpenStudy (anonymous):

wait so the common difference would be written 4,6,8,10?

OpenStudy (asnaseer):

look at the 4th term, it is simply: 2+4+6+8 do you know the formulae for the SUM of an arithmetic sequence?

OpenStudy (asnaseer):

\[S_n=\frac{n}{2}(2a_1+(n-1)d)\]

OpenStudy (asnaseer):

in your case:\[n=4, a_1=2, d=2\]

OpenStudy (asnaseer):

similarly, the 3rd term is just the SUM of the 1st 3 terms the 2nd term is the SUM of the 1st 2 terms so the nth term would be just the SUM of the 1st n terms

OpenStudy (anonymous):

so what would be the rule , how do i find that

OpenStudy (asnaseer):

I gave you the formulae above, it is:\[S_n=\frac{n}{2}(2a_1+(n-1)d)\] where, for your particular sequence:\[a_1=2, d=2\]

OpenStudy (asnaseer):

so for your sequence the nth term is given by:\[a_n=\frac{n}{2}(2*2+(n-1)*2)\]\[=\frac{n}{2}(4+2n-2)=\frac{n}{2}(2+2n)=n(n+1)\] do you see how we got to this?

OpenStudy (anonymous):

yeah i see now

OpenStudy (asnaseer):

these can be tricky, you just need to be patient and persevere...